# Holomorphic Deformations of Hyperbolicity Notions on Compact Complex Manifolds

### v1

@@ A Hermitian metric $\omega$ on $X$ is said to be :
where $\beta(z)=\lvert z\rvert^{2}=r^{2}$, hence $d\beta=2rdr$. Therefore, by combining (3.17) together (3.18), we get for any $r>r_{0}$ sufficiently big :
@@
$\overset{(d)}{\geq}\dfrac{2}{C_{0}B^{2}}\int_{0}^{r}\text{Vol}_{\omega,f}(\mathbb{B}_{t})dt=:CF(r)$
where $C:=\dfrac{2}{C_{0}B^{2}}>0$, and $F(r):=\int_{0}^{r}\text{Vol}_{\omega,f}(\mathbb{B}_{t})dt$. ($d$) follows from the condition (ii) in Definition 3.18. Thus, the previous inequality takes the form :
@@
Finally, we are willing to study the holomorphic deformations of $p$-HS hyperbolicity, $p$-Kähler hyperbolicity. Our first result in this direction is the deformation openness the of $p$-HS hyperbolicity property :
@@
Fix a $\mathscr{C}^{\infty}$-family $\{\omega_{t}\}_{t\in\mathbb{B}_{\varepsilon}(0)}$ of Hermitian metrics on $(X_{t})_{t\in\mathbb{B}_{\varepsilon}(0)}$. Let $\Omega$ be a $d$-closed $2p$-form on $X_{0}$ such that $\Omega^{p,p}$ is transverse and $\widetilde{\Omega}=d\Gamma$, for some $\widetilde{\omega_{0}}-$bounded (2p12p-1)-form on $\widetilde{X_{0}}$. For $t\in\mathbb{B}_{\varepsilon}(0)$, let $\Omega_{t}^{p,p}$ be the ($p,p$)-component of $\Omega$ with respect to the complex structure on $X_{t}$ (In particular : $\Omega_{0}^{p,p}=\Omega^{p,p}$, which is transverse by hypothesis). By continuity, $\Omega_{t}^{p,p}$ is transverse for $t$ close enough to 0 (see for instance Lemma 4.2, [8] or Proposition 4.12, [51] for the proof), meaning that $\Omega_{t}^{p,p}$ is a $p$-HS form on $X_{t}$, for $t\sim 0$. Morevoer, $\widetilde{\Omega}=d\Gamma$ with : $\lVert\Gamma\rVert_{L^{\infty}_{\widetilde{\omega_{t}}}(\widetilde{X_{t}})}\leq C_{t}\lVert\Gamma\rVert_{L^{\infty}_{\widetilde{\omega_{0}}}(\widetilde{X_{0}})}$$\infty$, for some $C_{t}>0$ (since the metrics $\omega_{0}$ and $\omega_{t}$ are comparable on the underlying compact manifold $X$). Hence $X_{t}$ is also $p$-HS hyperbolic for $t$ close enough to 0. ∎
@@
Let $\{X_{t}:t\in\mathbb{B}_{\varepsilon}(0)\}$ be an analytic family of compact complex $n$-dimensional manifolds such that $X_{0}$ is Kähler hyperbolic. Then the fibers $X_{t}$ are HS-hyperbolic for $t$ sufficiently small.
When $2\leq p\leq n-1$, the fibers are not necessarily $p$-Kähler. Sufficient conditions ensuring the stability of the $p$-Kähler property were given for instance in [50], Theorem 1.1 and [51], Theorem 1.8. In order to study the small deformations of $p$-Kähler hyperbolic manifolds, we will assume that the $\partial\bar{\partial}$-Lemma holds. Note that a $p$-Kähler $\partial\bar{\partial}$-manifold with $2\leq p\leq n-1$ is not necessarily Kähler (take a proper modification of a compact Kähler manifold, see Theorem 4.8, [5] and [12], Theorem 5.22). First, we recall the definition of Bott-Chern cohomology and Aeppli cohomology :
Definition 3.26 ([52], 2.b).
Let $X$ be a compact complex manifold with $\dim_{\mathbb{C}}X=n$. The $(p,q)$-th Bott-Chern cohomology of $X$ is defined as follows :
$H_{BC}^{p,q}(X,\mathbb{C})=\dfrac{\text{Ker}(\partial:\mathscr{C}^{\infty}_{p,q}(X)\longrightarrow\mathscr{C}^{\infty}_{p+1,q}(X))\cap\text{Ker}(\bar{\partial}:\mathscr{C}^{\infty}_{p,q}(X)\longrightarrow\mathscr{C}^{\infty}_{p,q+1}(X))}{\text{Im}(\partial\bar{\partial}:\mathscr{C}^{\infty}_{p-1,q-1}(X)\longrightarrow\mathscr{C}^{\infty}_{p,q}(X))}$
Note that the $L^{2}$-orthogonal decomposition (2.5) induced by the Bott-Chern Laplacian yields the Hodge isomorphism : $H_{BC}^{p,q}(X,\mathbb{C})\simeq\mathcal{H}_{\Delta_{BC}}^{p,q}(X)$, for any $p,q=0,\dots,n$.
Definition 3.27 ([52], 2.c).
Let $X$ be a compact complex manifold with $\dim_{\mathbb{C}}X=n$. The $(p,q)$-th Aeppli cohomology of $X$ is defined as follows :
$H_{A}^{p,q}(X,\mathbb{C})=\dfrac{\text{Ker}(\partial\bar{\partial}:\mathscr{C}^{\infty}_{p,q}(X)\longrightarrow\mathscr{C}^{\infty}_{p+1,q+1}(X))}{\text{Im}(\partial:\mathscr{C}^{\infty}_{p-1,q}(X)\longrightarrow\mathscr{C}^{\infty}_{p,q}(X))+\text{Im}(\bar{\partial}:\mathscr{C}^{\infty}_{p,q-1}(X)\longrightarrow\mathscr{C}^{\infty}_{p,q}(X))}$
Prior to examining the deformation theory of $p$-Kähler hyperbolic manifolds, we recall some of the key characterizations of $\partial\bar{\partial}$-manifolds, highlighting their favourable Hodge-theoretic properties (which is why they are sometimes referred to as cohomologically Kähler), following ideas from [12], [45] and [6] (see also the exposition in [49], §1.3).
Proposition 3.28.
  1. 1.
[…]
  2. 2.
  3. 3.
    $\forall p,q=0,\dots,n$, the canonical linear maps :
    are isomorphisms.
We are now prepared to state our deformation theorem for $p$-Kähler hyperbolic manifolds. This result extends part (1) of Theorem 2.17 as well as Theorem 4.10 of [31] in the $L^{\infty}$-setting, for arbitrary degree $1\leq p\leq n-1$:
@@
Let $\{\omega_{t}\}_{t\in\mathbb{B}_{\varepsilon}(0)}$ be a $\mathscr{C}^{\infty}$-family of Hermitian metrics on the fibers $X_{t}$, and let $\Omega$ be a $p$-Kähler hyperbolic form on $X_{0}$, that is $\widetilde{\Omega}=d\eta$ for some $\widetilde{\omega_{0}}-$bounded (2p12p-1)-form $\eta$ on $\widetilde{X_{0}}$. Thanks to the deformation openness of the $\partial\bar{\partial}$-property, The Hodge decomposition remains valid on the fibers $X_{t}$ (Proposition 3.28.2). Hence, using Proposition 3.28.3, $\Omega$ decomposes with respect to the complex structure on $X_{t}$ for each $t\in\mathbb{B}_{\varepsilon}(0)$ as :
$\Omega=\underset{j+k=2p}{\sum}F_{BC,t}^{k,j}\Omega_{t}^{k,j}+du_{t}\hskip 2.0pt,\hskip 23.49976ptu_{t}\in\mathscr{C}_{2p-1}^{\infty}(X_{t})$ (3.19)
@@
where $F_{BC,t}^{k,j}:\mathscr{C}^{\infty}_{k,j}(X_{t})\longrightarrow\mathcal{H}_{BC}^{k,j}(X_{t})$ is the projection of the ($k,j$)-th component of $\Omega$ onto the Bott-Chern harmonic ($k,j$)-forms, for $k+j=2p$. Since $X_{0}$ ia a $\partial\bar{\partial}$-manifold, the function $t\in\mathbb{B}_{\varepsilon}(0)\longmapsto h^{p,p}_{BC}(X_{t})$ is constant (after shrinking $\varepsilon$ if necessary, see [6], Corollary 3.7). Then one defines :
$\Omega_{t}:=(F_{BC,t}+\partial_{t}\bar{\partial}_{t}\bar{\partial}_{t}^{*}\partial_{t}^{*}\mathbb{G}_{BC,t})\Omega_{t}^{p,p}\hskip 2.0pt,\hskip 23.49976pt\forall t\in\mathbb{B}_{\varepsilon}(0)$ (3.20)
@@
where $\mathbb{G}_{BC,t}:\mathscr{C}_{p,p}^{\infty}(X_{t})\longrightarrow\mathscr{C}_{p,p}^{\infty}(X_{t})$ is the Green operator associated to the Bot-Chern Laplacian $\Delta_{BC,t}$. $\Omega_{t}$ defines a $\mathscr{C}^{\infty}$-family of $d$-closed ($p,p$)-forms (thanks to Theorem 5 in [32]) with :
$\Omega_{0}=(F_{BC}+\partial\bar{\partial}\bar{\partial}^{*}\partial^{*}\mathbb{G}_{BC})\Omega_{0}^{p,p}\overset{(*)}{=}F_{BC}\Omega+\Delta_{BC}\mathbb{G}_{BC}\Omega=\Omega$
(where ()(*) follows from the fact that $\Omega$ is $d$-closed). Hence $\Omega_{t}$ is a $p$-Kähler form on $X_{t}$ for $t\sim 0$. By lifting (3.19) to the universal cover and taking the ($p,p$)-component we get :
$\widetilde{F_{BC,t}^{p,p}\Omega_{t}^{p,p}}=\partial_{t}(\eta_{t}^{p-1,p}-\pi^{*}u_{t}^{p,p-1})+\bar{\partial}_{t}(\eta_{t}^{p,p-1}-\pi^{*}u_{t}^{p,p-1})\hskip 2.0pt,\hskip 23.49976ptt\sim 0$ (3.21)
@@
$\widetilde{\Omega_{t}}=\partial_{t}(\eta_{t}^{p-1,p}+\pi^{*}(\bar{\partial}_{t}\bar{\partial}_{t}^{*}\partial_{t}^{*}\mathbb{G}_{BC,t}\Omega_{t}^{p,p}-u_{t}^{p,p-1})+\bar{\partial}_{t}(\eta_{t}^{p,p-1}-\pi^{*}u_{t}^{p,p-1})=\partial_{t}\alpha_{t}+\bar{\partial}_{t}\beta_{t}$ (3.22)
where $\alpha_{t}:=\eta_{t}^{p-1,p}+\pi^{*}(\bar{\partial}_{t}\bar{\partial}_{t}^{*}\partial_{t}^{*}\mathbb{G}_{BC,t}\Omega_{t}^{p,p}-u_{t}^{p,p-1})$ and $\beta_{t}=\eta_{t}^{p,p-1}-\pi^{*}u_{t}^{p,p-1}$, with :
$\begin{cases}\lVert\alpha_{t}\rVert_{L^{\infty}_{\widetilde{\omega_{t}}}(\widetilde{X}_{t})}&\leq\lVert\eta_{t}^{p-1,p}\rVert_{L^{\infty}_{\widetilde{\omega_{t}}}(\widetilde{X}_{t})}+\lVert\pi^{*}(\bar{\partial}_{t}\bar{\partial}_{t}^{*}\partial_{t}^{*}\mathbb{G}_{BC,t}\Omega_{t}^{p,p}-u_{t}^{p,p-1})\rVert_{L^{\infty}_{\widetilde{\omega_{t}}}(\widetilde{X}_{t})}\\ &\leq C_{t}\lVert\eta\rVert_{L^{\infty}_{\widetilde{\omega_{0}}}(\widetilde{X})}+\lVert\bar{\partial}_{t}\bar{\partial}_{t}^{*}\partial_{t}^{*}\mathbb{G}_{BC,t}\Omega_{t}^{p,p}\rVert_{L^{\infty}_{\omega_{t}}(X_{t})}+\lVert u_{t}\rVert_{L^{\infty}_{\omega_{t}}(X_{t})}<\infty\\ \lVert\beta_{t}\rVert_{L^{\infty}_{\widetilde{\omega_{t}}}(\widetilde{X}_{t})}&\leq\lVert\eta_{t}^{p,p-1}\rVert_{L^{\infty}_{\widetilde{\omega_{t}}}(\widetilde{X}_{t})}+\lVert\pi^{*}u_{t}^{p,p-1}\rVert_{L^{\infty}_{\widetilde{\omega_{t}}}(\widetilde{X}_{t})}\leq C_{t}\lVert\eta\rVert_{L^{\infty}_{\widetilde{\omega_{0}}}(\widetilde{X})}+\lVert u_{t}\rVert_{L^{\infty}_{\omega_{t}}(X)}<\infty\end{cases}$
@@
The $\partial\bar{\partial}$-assumption appears to be quite strong. Since, as mentioned above, D. Popovici conjectured that every $\partial\bar{\partial}$-manifold is balanced, and the Fino-Vezzoni conjecture predicts that any balanced SKT manifold is in fact Kähler ([16], Problem 3). This suggests a deeper relationship between these structures. Motivated by this, we propose the following conjecture in degree $1\leq p\leq n-1$ :
@@
An affirmative answer to this question would imply the deformation openness of $p$-SKT hyperbolicity, provided the $\partial\bar{\partial}$-property holds.

### v4

@@ 3.2  $p$-SKT and $p$-HS Hyperbolicity
where $\beta(z)=\lvert z\rvert^{2}=r^{2}$, hence $d\beta=2rdr$. Therefore, by combining (3.16) and (3.17), we get for any $r>r_{0}$ sufficiently large:
@@
$\overset{(c)}{\geq}\dfrac{2}{C_{0}B^{2}}\int_{0}^{r}\text{Vol}_{\omega,f}(\mathbb{B}_{t})dt=:CF(r)$
where $C:=\dfrac{2}{C_{0}B^{2}}>0$, $F(r):=\int_{0}^{r}\text{Vol}_{\omega,f}(\mathbb{B}_{t})dt$ and ($c$) follows from the condition (ii) in Definition 3.20. Thus, the previous inequality takes the form:
@@
3.3  Deformation Stability Results
Finally, we will study the holomorphic deformations of $p$-HS hyperbolic and $p$-Kähler hyperbolic manifolds. Our first result in this direction is the deformation openness of the $p$-HS hyperbolicity property:
@@ 3.3  Deformation Stability Results
Fix a $\mathscr{C}^{\infty}$-family $\{\omega_{t}\}_{t\in\mathbb{B}_{\varepsilon}(0)}$ of Hermitian metrics on $(X_{t})_{t\in\mathbb{B}_{\varepsilon}(0)}$. Let $\Omega$ be a $d$-closed $2p$-form on $X_{0}$ such that $\Omega^{p,p}$ is weakly strictly positive and $\widetilde{\Omega}=d\Gamma$, for some $\widetilde{\omega_{0}}-$bounded (2p12p-1)-form on $\widetilde{X_{0}}$. For $t\in\mathbb{B}_{\varepsilon}(0)$, let $\Omega_{t}^{p,p}$ be the ($p,p$)-component of $\Omega$ with respect to the complex structure on $X_{t}$ (In particular: $\Omega_{0}^{p,p}=\Omega^{p,p}$, which is weakly strictly positive by hypothesis). By continuity, $\Omega_{t}^{p,p}$ is weakly strictly positive for $t$ close enough to 0 (see for instance Lemma 4.2, [8] or Proposition 4.12, [59] for the proof), meaning that $\Omega_{t}^{p,p}$ is a $p$-HS form on $X_{t}$, for $t\sim 0$. Morevoer, $\widetilde{\Omega}=d\Gamma$ with: $\lVert\Gamma\rVert_{L^{\infty}_{\widetilde{\omega_{t}}}(\widetilde{X_{t}})}\leq C_{t}\lVert\Gamma\rVert_{L^{\infty}_{\widetilde{\omega_{0}}}(\widetilde{X_{0}})}$$\infty$, for some $C_{t}>0$ (since the metrics $\omega_{0}$ and $\omega_{t}$ are comparable on the underlying compact manifold $X$). Hence $X_{t}$ is also $p$-HS hyperbolic for $t$ close enough to 0. ∎
@@
Let $\{X_{t}:t\in\mathbb{B}_{\varepsilon}(0)\}$ be an analytic family of compact complex manifolds of dimension $n$, such that $X_{0}$ is Kähler hyperbolic. Then the fibers $X_{t}$ are HS-hyperbolic for $t$ sufficiently small.
When $2\leq p\leq n-1$, the fibers are not necessarily $p$-Kähler. Sufficient conditions ensuring the stability of the $p$-Kähler property were given for instance in [58], Theorem 1.1 and [59], Theorem 1.8. In order to study the small deformations of $p$-Kähler hyperbolic manifolds, we will assume that the $\partial\bar{\partial}$-property is satisfied. Note that a $p$-Kähler $\partial\bar{\partial}$-manifold with $2\leq p\leq n-1$ is not necessarily Kähler (take a proper modification of a compact Kähler manifold, see Theorem 4.8 in [5] and Theorem 5.22 in [13]). The following result extends part (1) of Theorem 2.15 as well as Theorem 4.10 of [33] in the $L^{\infty}$-setting, for arbitrary degree $1\leq p\leq n-1$:
@@
Let $\{\omega_{t}\}_{t\in\mathbb{B}_{\varepsilon}(0)}$ be a $\mathscr{C}^{\infty}$-family of Hermitian metrics on the fibers $X_{t}$, and let $\Omega$ be a $p$-Kähler hyperbolic form on $X_{0}$, that is $\widetilde{\Omega}=d\eta$ for some $\widetilde{\omega_{0}}-$bounded (2p12p-1)-form $\eta$ on $\widetilde{X_{0}}$. Recall that the Bott-Chern Laplacian $\Delta_{BC}$ (with respect to the metric $\omega_{0}$), being elliptic and formally self-adjoint, induces the following $L^{2}_{\omega_{0}}$-orthogonal three-space decomposition in any bidegree ($p,q$) (see [53], §2):
$\mathscr{C}^{\infty}_{p,q}(X_{0})=\mathcal{H}_{BC}^{p,q}(X_{0})\oplus\text{Im}\hskip 2.0pt\partial\bar{\partial}\oplus\left(\text{Im}\hskip 2.0pt\partial^{*}+\text{Im}\hskip 2.0pt\bar{\partial}^{*}\right)\qquad,$ (3.18)
and we have
$\text{Ker}\hskip 2.0pt\bar{\partial}\cap\text{Ker}\hskip 2.0pt\partial=\mathcal{H}_{BC}^{p,q}(X_{0})\oplus\text{Im}\hskip 2.0pt\partial\bar{\partial}$ (3.19)
yielding the Hodge isomorphism $\mathcal{H}_{BC}^{p,q}(X_{0})\simeq H_{BC}^{p,q}(X_{0})$. Now, since $\Omega$ is real (as $\Omega$ is weakly strongly positive), $d\Omega=0$ implies that $\partial\Omega=\bar{\partial}\Omega=0$. Hence $\Omega\in\mathcal{H}_{BC}^{p,p}(X_{0})\oplus\text{Im}\hskip 2.0pt\partial\bar{\partial}$ thanks to (3.19). Moreover, $\Omega=F^{p,p}_{BC}\Omega+\Delta_{BC}\mathbb{G}_{BC}\Omega$ by definition of the Green operator $\mathbb{G}_{BC}$, where $F^{p,p}_{BC}\Omega$ is the projection of $\Omega$ onto the Bott-Chern harmonic space $\mathcal{H}^{p,p}_{BC}(X_{0})$.
The uniqueness of the decompostion of $\Omega$ with respect to (3.18) yields: $\Omega=F^{p,p}_{BC}\Omega+\partial\bar{\partial}\alpha$, for some $\alpha\in\mathscr{C}^{\infty}_{p-1,p-1}(X_{0})$. We can take $\alpha$ to be the minimal $L^{2}_{\omega_{0}}$-norm solution of this equation, which is given by the following Neumann-type formula ([53], Theorem 4.1):
$\alpha=(\partial\bar{\partial})^{*}\mathbb{G}_{BC}(\Omega-F_{BC}^{p,p}\Omega)=(\partial\bar{\partial})^{*}\mathbb{G}_{BC}\Omega$
@@
In other words, one can write $\Omega$ as: $\Omega=(F_{BC}^{p,p}+\partial\bar{\partial}\bar{\partial}^{*}\partial^{*}\mathbb{G}_{BC})\Omega$. We decompose $\Omega$ as: $\Omega=\underset{j+k=2p}{\sum}\Omega_{t}^{k,j}$. Since $X_{0}$ ia a $\partial\bar{\partial}$-manifold, the function $t\in\mathbb{B}_{\varepsilon}(0)\longmapsto h^{p,p}_{BC}(X_{t})$ is constant (after shrinking $\varepsilon$ if necessary, see [6], Corollary 3.7). Then one defines:
$\Omega_{t}:=(F_{BC,t}^{p,p}+\partial_{t}\bar{\partial}_{t}\bar{\partial}_{t}^{*}\partial_{t}^{*}\mathbb{G}_{BC,t})\Omega_{t}^{p,p}\hskip 2.0pt,\hskip 23.49976pt\forall t\in\mathbb{B}_{\varepsilon}(0)$ (3.20)
@@
where $\mathbb{G}_{BC,t}:\mathscr{C}_{p,p}^{\infty}(X_{t})\longrightarrow\mathscr{C}_{p,p}^{\infty}(X_{t})$ is the Green operator associated with the Bot-Chern Laplacian $\Delta_{BC,t}$, and $F_{BC,t}^{p,p}:\mathscr{C}^{\infty}_{p,p}(X_{t})\longrightarrow\mathcal{H}_{BC}^{p,p}(X_{t})$ is the projection onto the Bott-Chern harmonic ($p,p$)-forms. $\{\Omega_{t}\}_{t\in\mathbb{B}_{\varepsilon}(0)}$ defines a $\mathscr{C}^{\infty}$-family of $d$-closed ($p,p$)-forms (thanks to Theorem 5 in [35]) with $\Omega_{0}=\Omega$. In particular, $\{\Omega_{t}\}_{t\sim 0}$ is a $\mathscr{C}^{\infty}$-family of $p$-Kähler forms on the respective fibers $X_{t}$ for $t\sim 0$, with $\Omega_{0}=\Omega$.
Now, $d\Omega=0$ implies: $\partial_{t}\bar{\partial}_{t}\Omega_{t}^{k,j}=0$, with $k+j=2p$, $\forall t\sim 0$. Using the fact that the canonical isomorphism given in Proposition 2.8.(3) is induced by the identity map (which exists thanks to the deformation openness of the $\partial\bar{\partial}$-property), we get:
$\Omega_{t}^{k,j}]_{A}=[F_{BC,t}^{k,j}\Omega_{t}^{k,j}]_{A}\qquad,\hskip 23.49976pt\text{ with }k=j=2p\qquad,\qquad\forall t\sim 0.$
as $\Omega_{t}^{k,j}]_{BC}=[F_{BC,t}^{k,j}\Omega_{t}^{k,j}]_{BC}$. This implies that: $\Omega_{t}^{k,j}=F_{BC,t}^{k,j}\Omega_{t}^{k,j}+\partial_{t}u_{t}^{k-1,j}+\bar{\partial}_{t}u_{t}^{k,j-1}$, with $k+j=2p$, $t\sim 0$, and $u_{t}^{k-1,j}\in\mathscr{C}^{\infty}_{k-1,j}(X_{t})$, $u_{t}^{k,j-1}\in\mathscr{C}^{\infty}_{k,j-1}(X_{t})$.
Hence $\Omega$ decomposes with respect to the complex structure on $X_{t}$ for each $t\in\mathbb{B}_{\varepsilon}(0)$ as:
$\Omega=\underset{j+k=2p}{\sum}F_{BC,t}^{k,j}\Omega_{t}^{k,j}+\partial_{t}v_{t}+\bar{\partial}_{t}w_{t}\hskip 2.0pt\qquad,\hskip 23.49976ptv_{t},w_{t}\in\mathscr{C}_{2p-1}^{\infty}(X_{t})\qquad,\qquad\forall t\sim 0,$ (3.21)
where $F_{BC,t}^{k,j}:\mathscr{C}^{\infty}_{k,j}(X_{t})\longrightarrow\mathcal{H}_{BC}^{k,j}(X_{t})$ is the projection onto the Bott-Chern harmonic ($k,j$)-forms. By lifting (3.21) to the universal cover and taking the ($p,p$)-component we get:
$\widetilde{F_{BC,t}^{p,p}\Omega_{t}^{p,p}}=\partial_{t}(\eta_{t}^{p-1,p}-\pi^{*}v_{t}^{p-1,p})+\bar{\partial}_{t}(\eta_{t}^{p,p-1}-\pi^{*}w_{t}^{p,p-1})\hskip 2.0pt,\hskip 23.49976ptt\sim 0.$ (3.22)
@@
$\widetilde{\Omega_{t}}=\partial_{t}(\eta_{t}^{p-1,p}+\pi^{*}(\bar{\partial}_{t}\bar{\partial}_{t}^{*}\partial_{t}^{*}\mathbb{G}_{BC,t}\Omega_{t}^{p,p}-v_{t}^{p-1,p})+\bar{\partial}_{t}(\eta_{t}^{p,p-1}-\pi^{*}w_{t}^{p,p-1})=\partial_{t}\alpha_{t}+\bar{\partial}_{t}\beta_{t}$ (3.23)
where $\alpha_{t}:=\eta_{t}^{p-1,p}+\pi^{*}(\bar{\partial}_{t}\bar{\partial}_{t}^{*}\partial_{t}^{*}\mathbb{G}_{BC,t}\Omega_{t}^{p,p}-v_{t}^{p-1,p})$ and $\beta_{t}=\eta_{t}^{p,p-1}-\pi^{*}w_{t}^{p,p-1}$, with:
$\begin{cases}\lVert\alpha_{t}\rVert_{L^{\infty}_{\widetilde{\omega_{t}}}(\widetilde{X}_{t})}&\leq\lVert\eta_{t}^{p-1,p}\rVert_{L^{\infty}_{\widetilde{\omega_{t}}}(\widetilde{X}_{t})}+\lVert\pi^{*}(\bar{\partial}_{t}\bar{\partial}_{t}^{*}\partial_{t}^{*}\mathbb{G}_{BC,t}\Omega_{t}^{p,p}-v_{t}^{p-1,p})\rVert_{L^{\infty}_{\widetilde{\omega_{t}}}(\widetilde{X}_{t})}\\ &\leq C_{t}\lVert\eta\rVert_{L^{\infty}_{\widetilde{\omega_{0}}}(\widetilde{X})}+\lVert\bar{\partial}_{t}\bar{\partial}_{t}^{*}\partial_{t}^{*}\mathbb{G}_{BC,t}\Omega_{t}^{p,p}\rVert_{L^{\infty}_{\omega_{t}}(X_{t})}+\lVert v_{t}\rVert_{L^{\infty}_{\omega_{t}}(X_{t})}<\infty\\ \lVert\beta_{t}\rVert_{L^{\infty}_{\widetilde{\omega_{t}}}(\widetilde{X}_{t})}&\leq\lVert\eta_{t}^{p,p-1}\rVert_{L^{\infty}_{\widetilde{\omega_{t}}}(\widetilde{X}_{t})}+\lVert\pi^{*}w_{t}^{p,p-1}\rVert_{L^{\infty}_{\widetilde{\omega_{t}}}(\widetilde{X}_{t})}\leq C_{t}\lVert\eta\rVert_{L^{\infty}_{\widetilde{\omega_{0}}}(\widetilde{X})}+\lVert w_{t}\rVert_{L^{\infty}_{\omega_{t}}(X)}<\infty\end{cases}$
@@
The $\partial\bar{\partial}$-assumption appears to be quite strong, since, as mentioned above, D. Popovici conjectured that every $\partial\bar{\partial}$-manifold is balanced, and the Fino-Vezzoni conjecture predicts that any balanced SKT manifold is in fact Kähler ([17], Problem 3). This suggests a deeper relationship between these structures. Motivated by this, we propose the following conjecture for any $1\leq p\leq n-1$:
@@
An affirmative answer to this question would imply the deformation openness of $p$-SKT hyperbolicity, provided the $\partial\bar{\partial}$-property holds.thereby answering positively Question 5.6 in [52]. It follows from Theorem 3.2.17.(2)
Acknowledgements
This work is part of the author’s PhD thesis, and he would like to thank his supervisor Dan Popovici for carefully reading the manuscript and for providing several very useful remarks. The author would also like to express his gratitude to Philippe Eyssidieux for pointing out the paper of Jean-Paul Mohsen on the construction of negatively curved complete intersections [46]. Many thanks are also due to Jeffrey Streets for drawing his attention to his very interesting joint work with Man-Chun Lee [37].